### v1

@@ Abstract
Precision and Recall are foundational metrics in machine learning applications where both accurate predictions and comprehensive coverage are essential, such as in recommender systems and multi-label learning. In these tasks, balancing precision (the proportion of relevant items among those predicted) and recall (the proportion of relevant items successfully predicted) is crucial. A key challenge is that one-sided feedback—where only positive examples are observed during training—is inherent in many practical problems. For instance, in recommender systems like YouTube, training data only consists of songs or videos that a user has actively selected, while unselected items remain unseen. Despite this lack of negative feedback in training, it becomes crucial at test time; for example, in a recommender system, it’s essential to avoid recommending items the user would likely dislike.
We introduce a Probably Approximately Correct (PAC) learning framework where each hypothesis is represented by a graph, with edges indicating positive interactions, such as between users and items. This framework subsumes the classical binary and multi-class PAC learning models as well as multi-label learning with partial feedback, where only a single random correct label per example is observed, rather than all correct labels.
Our work uncovers a rich statistical and algorithmic landscape, with nuanced boundaries on what can and cannot be learned. Notably, classical methods like Empirical Risk Minimization fail in this setting, even for simple hypothesis classes with only two hypotheses. To address these challenges, we develop novel algorithms that learn exclusively from positive data, effectively minimizing both precision and recall losses. Specifically, in the realizable setting, we design algorithms that achieve optimal sample complexity guarantees. In the agnostic case, we show that it is impossible to achieve additive error guarantees (i.e., additive regret)—as is standard in PAC learning—and instead obtain meaningful multiplicative approximations.
@@ 1 Introduction
Precision and Recall are fundamental metrics in a variety of machine learning applications where accurate prediction and comprehensive coverage are both critical. Recommender systems are one example of machine learning applications where precision and recall are essential. These systems, used in online platforms like streaming services and e-commerce websites, suggest items—such as movies, music, or products—that match user preferences. For instance, Netflix aims to recommend shows that a user will watch and enjoy, based on limited past interactions. Another example is multi-label learning, in which each input corresponds to multiple labels (for instance, an image may contain multiple objects), and the learning objective is to return all labels associated with each input.
@@
  • Precision- The proportion of recommended items that the user likes. Low precision loss means most suggested items are liked by the users.
  • Recall- The proportion of items the user would like that are successfully recommended. Low recall loss ensures the system does not miss out on suggesting items the user would have liked.
Precision and recall are often at odds. Increasing the number of recommendations can improve recall but may reduce precision. Beyond recommender systems, which will be our running example, these metrics play a vital role in multi-label learning, such as predicting all the objects present in an image, and in platforms like dating apps, where matching users effectively depends on balancing the relevance and variety of suggestions.
@@
In an ideal scenario, we might consider a full information model where, for each user in the training set, the algorithm has access to all the items they like. This setup implicitly includes negative examples, as items not on a user’s list are considered unliked. At test time, the algorithm would then predict a list of recommendations for a given random user. Such a setting aligns well with the standard Probably Approximately Correct (PAC) framework [Valiant84], allowing us to apply standard PAC solutions (e.g., Empirical Risk Minimization).
@@
However, this assumption of full information is often unrealistic in many applications. In real-world scenarios, we typically only observe a small fraction of the items a user liked in the training set, with no explicit information on items the user would not chosen. This setting is better characterized as a partial-information model. For instance, in Spotify, for each observed user in our training data, we might only know the songs they have listened to, without knowing about the vast majority they have not listened to. At test time, given a random user drawn from the same distribution as the training set, we present a full list of recommended items (and not just a single song).
Following standard practice in learning theory, we consider a hypothesis class $\mathcal{H}$. Each hypothesis in $\mathcal{H}$ is modeled by a finite graph, $g$. For example, in the context of recommender systems, nodes represent users and items, and an edge between a user and an item indicates that the item is liked by the user (and should therefore be recommended). Our goal is to return a graph that minimizes both precision and recall losses. The recall and precision losses of graph $g$ are measured with respect to the target graph $g^{target}$, which captures the true set of items liked by each user:
@@
$\ell^{\textrm{precision}}(g):=\mathbb{E}_{x\sim\mathcal{D}}\left[\frac{|N_{g}(x)\setminus N_{g^{\text{target}}}(x)|}{n_{g}(x)}\right]\,,$
$\ell^{\textrm{recall}}(g):=\mathbb{E}_{x\sim\mathcal{D}}\left[\frac{|N_{g^{\text{target}}}(x)\setminus N_{g}(x)|}{n_{g^{\text{target}}}(x)}\right]\,,$
where $N_{g^{\prime}}(x)$ denotes the neighborhood of a node $x$ in a graph $g^{\prime}$ and $n_{g^{\prime}}(x)=|N_{g^{\prime}}(x)|$ denotes the number of neighbors of $x$.
Note that binary and multiclass PAC learning can be viewed as special cases of our model, where the graphs are bipartite: one side contains nodes representing the inputs $x$, the other side contains nodes represent the labels $y$, and an edge $(x,y)$ indicates that $y$ is the label of $x$. We can also model multi-label PAC learning by allowing multiple edges $(x,y)$ where each $y$ represents one of the (potentially, multiple) labels of $x$.
@@
We distinguish between two settings- realizable and agnostic. In the realizable setting, we assume $g^{\text{target}}$ is in the class $\mathcal{H}$, and, aim to find an hypothesis with small precision and recall losses. In the agnostic setting, we do not assume that a perfect hypothesis is in the class. Instead, we aim to compete with the “best” hypothesis (graph) in the class, acknowledging that some error is unavoidable. In the context of the agnostic setting, defining the “best” hypothesis is subtle. One hypothesis might have high precision but low recall, while another has the opposite. Depending on the application’s needs, one may prefer higher precision over recall or vice versa.
@@
This naturally leads us to the concept of Pareto-loss objective, which captures the trade-offs between precision and recall along the Pareto frontier. [^1] The term “Pareto” originates from Vilfredo Pareto, an economist who observed that certain distributions followed a pattern where improvements in one dimension often involved trade-offs in another. The Pareto frontier, inspired by this principle, represents optimal trade-offs between competing objectives. Here, our Pareto loss captures the balance between precision and recall, aiming to improve both while acknowledging the inherent trade-off. Namely, the “best” graphs are on a Pareto frontier, which is the set of graphs where no other graphs are better in both precision and recall simultaneously (see, e.g., Figure 1). Here, we try, given desired precision and recall losses parameters $(p,r)$ to return a graph whose precision and recall losses $(p^{\prime},r^{\prime})$ satisfy $p^{\prime}\lesssim p$ and $r^{\prime}\lesssim r$. For example, one might aim to optimize precision while keeping recall below a specific threshold (e.g., at most $0.5$).
@@
Figure 1: Example of graphs with varying precision and recall losses. Each point is a distinct graph, with red points on the Pareto frontier, showing optimal trade-offs between precision and recall losses. The empty graph always achieves zero precision loss (but has no guarantee on the recall loss), while the complete graph always achieves zero recall loss (but has no guarantee on the precision loss).
@@
Our goal is to design algorithms whose sample complexity is polynomial in the log size of the class and the inverses of the accuracy and confidence parameters, and in some cases, on the maximum degree of the graph (which is arguably small in certain applications). We focus on finite hypothesis classes and aim for sample complexity bounds that depend logarithmically, rather than linearly, on the size of the class. This goal is motivated by standard sample complexity bounds in PAC learning and is particularly relevant when training data is costly to obtain, as is often the case with human-provided feedback. [^2] While it would be interesting to explore infinite hypothesis classes and characterize them via a combinatorial measure in the style of VC dimension, the findings in [lechner2024inherent] suggest that such a dimension may not exist in this setting. We leave this as an intriguing open question, as addressing it falls beyond the scope of this work, which focuses on the finite case—a setting that is already challenging. However, we aim to avoid dependencies on the number of vertices or edges in the graph.
@@
Our learning problem is significantly more challenging than standard supervised learning tasks due to the absence of negative examples. Namely, we only observe positive examples, making it impossible to estimate precision loss directly. Without knowing what items users dislike, we cannot estimate how many irrelevant items a hypothesis might recommend. This limitation challenges approaches like standard supervised learning, which rely on both positive and negative examples.
@@
In standard supervised learning, a classical solution known as Empirical Risk Minimization (ERM) involves finding a hypothesis that best fits the data by minimizing the average loss over all observed examples. However, in our case, the absence of negative examples means that ERM cannot be applied effectively, as there is no way to determine how well a hypothesis avoids irrelevant items. For example, the complete graph (where every possible recommendation is made) is consistent with every training set, since we have no negative examples to contradict it. However, such a hypothesis might have poor precision. Without negative examples in the training set, any hypothesis that covers all observed positive examples appears equally valid, even if it recommends many irrelevant items and incurs a high precision loss. The failure of the ERM principle is not unique to our learning problem; it also occurs in other learning problems, such as multi-class classification [daniely2015multiclass], density estimation [devroye2001combinatorial, bousquet2019optimal], and partial concept learning [alon2022theory].
@@
In fact, it is not just that ERM would fail; we provide an example in which two hypotheses have nearly the same recall loss but very different precision losses, making it impossible to determine which hypothesis has better precision loss based solely on the training data (see Example 1 for more details).
These challenges necessitate novel approaches to learning and evaluating hypotheses.
Contributions
@@
  • Learning Model: We propose a learning model that operates under partial information, where the algorithm observes only positive examples. For each user drawn from unknown distribution $\mathcal{D}$, it randomly observes one item they like. This model reflects real-world constraints and is more practical than assuming access to complete preference profiles as done in multi-label learning.
@@
    achieve recall and precision losses of at most $\varepsilon$ with probability at least $1-\delta$. We propose two distinct approaches to achieve this goal: the first circumvents the challenge of estimating precision by minimizing an appropriate surrogate loss, while the second takes a more intuitive approach inspired by the maximum likelihood principle [Shalev-Shwartz_Ben-David_2014]. In essence, this second algorithm, when presented with two graphs consistent with the data, prioritizes the one with smaller degrees in a suitably quantified sense.
  • Agnostic Setting: We demonstrate that achieving a vanishing additive error, as is standard in learning theory, is impossible in this setting by providing lower bounds on the sum of precision and recall losses, with multiplicative factors greater than $1$. In the other direction, we show that constant multiplicative factor guarantees are indeed achievable by adapting our realizable setting algorithms to the agnostic case. Closing the gap between our upper and lower bounds on the best achievable multiplicative factor ($5$ vs. $1.05$) remains an open question. For the Pareto-loss case, we establish both upper and lower bounds for the following question: Given that there exists a hypothesis in the class with precision and recall losses $(p,r)$, which guarantee pairs $(p^{\prime},r^{\prime})$ are achievable? Finally, we pose open questions about determining the optimal factors achievable in the agnostic setting.
@@
Precision and recall are natural and standard metrics used broadly in machine learning, spanning applications from binary classification [juba2019precision], multi-class classification [grandini2020metrics], regression [torgo2009precision], and time series [DBLP:conf/nips/TatbulLZAG18] to information retrieval [arora2016evaluation] and generative models [sajjadi2018assessing]. Beyond precision-recall, another related metric—the area under the ROC curve (AUC)—has also been extensively studied in the history of binary classification [DBLP:conf/nips/CortesM03, DBLP:conf/nips/CortesM04, DBLP:conf/icml/Rosset04, DBLP:journals/jmlr/AgarwalGHHR05], with a focus on generalization. Our work, however, studies a different problem of multi-label learning where the goal is to recommend a list of items to each user. Recommending a list of items has also been addressed in the context of cascading bandits [kveton2015cascading]. However, while our objective is to identify the items that each user likes, their focus is on learning the top $K$ items that are liked by most users. Another feature of our learning problem is that we can only learn from positive examples. PAC learning for binary classification from positive examples has been studied in the literature [denis1998pac, de1999positive, letouzey2000learning, bekker2020learning].
Multi-label learning [McCallum1999Multi, Schapire2000BoosTexterAB] has been an area of study in machine learning, with various, primarily experimental approaches (see, e.g., [Elisseeff01, Petterson11, NKapoor12] and [Zhang14survey, BOGATINOVSKI2022survey] for surveys). In multi-label learning, the training set consists of examples, each associated with multiple labels rather than just one. The goal is to train a model that can learn the relationships between the features of each example and all its labels. At test time, the learner predicts a list of labels for new examples, aiming to capture all the relevant labels that apply, rather than just a single one.
Some works have examined multi-label learning from a theoretical standpoint, focusing in particular on the Bayes consistency of surrogate losses. Bayes-consistency in multi-label learning ensures that minimizing a surrogate loss also leads to minimizing the true target loss, which is crucial in multi-label settings where optimizing the actual loss is often computationally infeasible as it is non-convex, discrete losses in multi-label settings. Initiated by  [pmlr-v19-gao11a] who first addressed Bayes-consistency for Hamming and ranking losses, showing binary relevance’s consistency with Hamming loss but highlighting ranking loss difficulties. Extensions include rank-based metrics like precision@$\kappa$ and recall@$\kappa$ [Menon19], which are loss functions defined under the constraint that the number of labels predicted by the model is limited to $\kappa$. Recently, [mao2024multilabellearningstrongerconsistency] established $H$-consistency bounds for multi-label learning, offering stronger guarantees than Bayes-consistency by providing non-asymptotic guarantees that apply to finite number of samples. Our model is inherently more challenging than traditional multi-label learning because our training set consists of examples, each associated with only a single correct label rather than all possible correct labels, with no negative feedback. Yet, at test time, the learner still needs to predict a list of relevant labels for new examples.

### v2

@@ Abstract
Precision and Recall are fundamental metrics in machine learning tasks where both accurate predictions and comprehensive coverage are essential, such as in multi-label learning, language generation, medical studies, and recommender systems. A key challenge in these settings is the prevalence of one-sided feedback, where only positive examples are observed during training—e.g., in multi-label tasks like tagging people in Facebook photos, we may observe only a few tagged individuals, without knowing who else appears in the image. To address learning under such partial feedback, we introduce a Probably Approximately Correct (PAC) framework in which hypotheses are set functions that map each input to a set of items, extending beyond single-label predictions and generalizing classical binary, multi-class, and multi-label models. Our results reveal sharp statistical and algorithmic separations from standard settings: classical methods such as Empirical Risk Minimization provably fail, even for simple hypothesis classes. We develop new algorithms that learn from positive data alone, achieving optimal sample complexity in the realizable case, and establishing multiplicative—rather than additive—approximation guarantees in the agnostic case, where achieving additive regret is impossible.
@@ 1 Introduction
Precision and Recall are fundamental metrics in many machine learning applications, including multi-label learning, medical studies, generative models, information retrieval, and recommender systems, where the goal is to learn a set for each input and both accurate prediction and comprehensive coverage are critical. For example, in multi-label learning—where each input may correspond to multiple labels (e.g., an image containing multiple objects or a group photo containing multiple people)—the objective is to return the complete set of labels associated with each input. Recommender systems also rely heavily on precision and recall; for instance, Netflix aims to recommend a list of shows that includes all those a user would enjoy and excludes those they would not. In medical studies, precision, called positive predictive value (PPV), and recall, called sensitivity, are key performance metrics. For instance, if a test predicts that $10$ patients have a condition and $9$ truly have it, its PPV (precision) is $90\%$. If $5$ patients actually have the condition and the test correctly identifies $4$ of them, its sensitivity (recall) is $80\%$. In language identification and generation in the limit [Gol67, KM24, KMV24], hallucination denotes a failure of precision, where the model produces strings outside the true language, whereas mode collapse denotes a failure of recall, where the model’s outputs become so constrained that valid, previously unseen strings are never generated .
@@
  • Precision — the proportion of output items that are in the ground-truth set. Low precision loss means that most output items are correct.
  • Recall — the proportion of ground-truth items that are included in the output. Low recall loss means the system captures most of the correct items.
@@
Precision and recall are often at odds. Increasing the size of output set can improve recall but may reduce precision. In an ideal scenario, we might consider a full information model where, for each input in the training data, the algorithm has access to the set of all the ground-truth items. This setup implicitly includes negative examples, as items not in the set are known to be incorrect. At test time, the algorithm would then predict a set of items for each given input. Such a setting aligns well with the standard Probably Approximately Correct (PAC) framework [Val84], allowing us to apply standard PAC solutions (e.g., Empirical Risk Minimization).
@@
However, the assumption of full information is often unrealistic in many applications. In real-world scenarios, we typically observe only a small fraction of the ground-truth items (e.g., some objects in an image) during training, with no explicit information about unobserved items (e.g., which objects are not in the image). This situation is better characterized as a partial-information model. For instance, on Facebook, for each photo in the training data, we might only know some of the tagged individuals, without knowing whether any untagged person appears in the photo or not. To formalize this setting, we consider a simple abstraction: during training, for each input, we observe only a single item sampled uniformly at random from its ground-truth set. At test time, given a random input, the model is expected to return the full set of associated items, not just one.
Following standard practice in learning theory, given an input space $\mathcal{X}$ and a label space $\mathcal{Y}$, we consider a hypothesis class $\mathcal{H}$. Each hypothesis is a set function $g:\mathcal{X}\to 2^{\mathcal{Y}}$ that maps each input in $\mathcal{X}$ to a set of labels in $\mathcal{Y}$. For example, in the context of multi-label learning, inputs may be images and items may be objects; a hypothesis maps an image to the set of objects it contains. Our goal is to find a set function that minimizes both precision and recall losses. These losses are defined with respect to a target hypothesis $g^{\text{target}}$, which captures the ground-truth label sets, using the standard counting measure:
@@
$\ell^{\textrm{precision}}(g)$ $:=\mathbb{E}_{x\sim\mathcal{D}}\left[\frac{|g(x)\setminus g^{\text{target}}(x)|}{|g(x)|}\right]\quad\mbox{and}\quad\ell^{\textrm{recall}}(g):=\mathbb{E}_{x\sim\mathcal{D}}\left[\frac{|g^{\text{target}}(x)\setminus g(x)|}{|g^{\text{target}}(x)|}\right]\,$,
where $\mathcal{D}$ denotes the distribution over inputs. Precision and recall are equal to $1$ minus the precision and recall losses, respectively. Binary and multiclass PAC learning can be viewed as special cases of our model, where we restricted the set size to be $1$. Note that, for each given $x$, we adopt the standard counting measure to quantify the difference between $g(x)$ and $g^{\text{target}}(x)$. While other measures are possible, it is often challenging to define more sophisticated alternatives in a principled and intuitive way. To the best of our knowledge, the only generalization of precision and recall beyond counting measure is proposed by [SBLBG18], though the resulting notions are relatively complex. Hence, we will be focusing on the counting measure in this paper and leave a more general notion of precision and recall as a open question.
@@
We distinguish between two settings- realizable and agnostic. In the realizable setting, we assume the target hypothesis $g^{\text{target}}$ is in the class $\mathcal{H}$, and, aim to find a hypothesis with small precision and recall losses. In the agnostic setting, we do not assume that a perfect hypothesis is in the class. Instead, we aim to compete with the “best” hypothesis in the class, acknowledging that some error is unavoidable. In the context of the agnostic setting, defining the “best” hypothesis is subtle. One hypothesis might have high precision but low recall, while another has the opposite. Depending on the application’s needs, one may prefer higher precision over recall or vice versa.
@@
This naturally leads us to the concept of Pareto-loss objective, which captures the trade-offs between precision and recall along the Pareto frontier. [^1] The term “Pareto” originates from Vilfredo Pareto, an economist who observed that certain distributions followed a pattern where improvements in one dimension often involved trade-offs in another. The concept of Pareto frontier, inspired by this principle, represents optimal trade-offs between competing objectives. Our Pareto loss captures the balance between precision and recall, aiming to improve both while acknowledging the trade-off.
Namely, the “best” hypotheses are on a Pareto frontier, which is the set of hypotheses where no other hypotheses are better in both precision and recall simultaneously (see, e.g., Figure 1). Here, we try, given desired precision and recall losses parameters $(p,r)$ to return a hypothesis whose precision and recall losses $(p^{\prime},r^{\prime})$ satisfy $p^{\prime}\lesssim p$ and $r^{\prime}\lesssim r$. For example, one might aim to optimize precision while keeping recall below a specific threshold (e.g., at most $0.5$).
@@
Figure 1: Example of hypotheses with varying precision and recall losses. Each point is a distinct hypothesis, with red points on the Pareto frontier, showing optimal trade-offs between precision and recall losses. The empty set function (which always return the empty set) always achieves zero precision loss (but has no guarantee on the recall loss), while the complete set function (which always return the entire label space $\mathcal{Y}$ for any given input $x$) always achieves zero recall loss (but has no guarantee on the precision loss).
@@
Our goal is to design algorithms whose sample complexity is polynomial in the log size of the class and the inverses of the accuracy and confidence parameters, and in some cases, on the maximum size of the output (which is arguably small in certain applications). We focus on finite hypothesis classes and aim for sample complexity bounds that depend logarithmically, rather than linearly, on the size of the class. This goal is motivated by standard sample complexity bounds in PAC learning and is particularly relevant when training data is costly to obtain, as is often the case with human-provided feedback. [^2] While it would be interesting to explore infinite hypothesis classes and characterize them via a combinatorial measure in the style of VC dimension, the findings in [LB24] suggest that such a dimension may not exist in this setting. We leave this as an intriguing open question, as addressing it falls beyond the scope of this work, which focuses on the finite case—a setting that is already challenging.
@@
Our learning problem is significantly more challenging than standard supervised learning tasks due to the absence of negative examples—that is, input-label pairs where the label is not in the ground-truth set. We observe only positive examples, making it impossible to estimate precision loss directly: without knowing which labels are incorrect, we cannot determine how many incorrect labels a hypothesis might output. This limitation undermines standard supervised learning approaches, which rely on access to both positive and negative examples.
@@
In standard supervised learning, a classical solution known as Empirical Risk Minimization (ERM) involves finding a hypothesis that best fits the data by minimizing the average loss over all observed examples. However, in our case, the absence of negative examples means that ERM cannot be applied effectively, as there is no way to determine how well a hypothesis avoids incorrect labels. For example, the complete set function (where $g(x)=\mathcal{Y}$ outputs all items) is consistent with every training set, since we have no negative examples to contradict it. However, such a hypothesis might have poor precision. Without negative examples in the training set, any hypothesis that covers all observed positive examples appears equally valid, even if it recommends many irrelevant items and incurs a high precision loss. The failure of the ERM principle is not unique to our learning problem; it also occurs in other learning problems, such as multi-class classification [DSBS15], density estimation [DL01, BKM19], and partial concept learning [AHHM22].
@@
In fact, it is not just that ERM would fail; we provide an example in which two hypotheses have nearly the same recall loss but very different precision losses, making it impossible to determine which hypothesis has better precision loss based solely on the training data (see Example 1 for more details). These challenges necessitate novel approaches to learning and evaluating hypotheses.
Our Contributions
@@
  • Learning Model: We propose a learning model that operates under partial information, where the algorithm observes only positive examples. For each input drawn from unknown distribution $\mathcal{D}$, it observes one label sampled uniformly at random from its ground-truth set. This model reflects real-world constraints and is more practical than assuming access to complete ground-truth set as done in multi-label learning.
@@
    achieve recall and precision losses of at most $\varepsilon$ with probability at least $1-\delta$. We propose two distinct approaches to achieve this goal: the first circumvents the challenge of estimating precision by minimizing an appropriate surrogate loss, while the second takes a more intuitive approach inspired by the maximum likelihood principle [SB14]. In essence, this second algorithm, when presented with two hypotheses consistent with the data, prioritizes the one with smaller output sizes in a suitably quantified sense.
  • Agnostic Setting: We demonstrate that achieving a vanishing additive error, as is standard in learning theory, is impossible in this setting by providing lower bounds on the sum of precision and recall losses, with multiplicative factors greater than $1$. In the other direction, we show that constant multiplicative factor guarantees are indeed achievable by adapting our realizable setting algorithms to the agnostic case. Closing the gap between our upper and lower bounds on the best achievable multiplicative factor ($5$ vs. $1.05$) remains an open question. For the Pareto-loss case, we establish both upper and lower bounds for the following question: Given that there exists a hypothesis in the class with precision and recall losses $(p,r)$, which guarantee pairs $(p^{\prime},r^{\prime})$ are achievable? Finally, we pose open questions about the optimal factors in the agnostic setting.
@@
Precision and recall are natural and standard metrics used broadly in machine learning, spanning applications from binary classification [JL19, DT21], multi-class classification [GBV20], regression [TR09], and time series [TLZAG18] to information retrieval [AKV16] and generative models [SBLBG18]. Beyond precision-recall, another related metric—the area under the ROC curve (AUC)—has also been extensively studied in the history of binary classification [CM03, CM04, Ros04, AGHHR05], with a focus on generalization. Our work, however, studies a different problem of multi-label learning where the goal is to output a list of labels for each input. In the context of recommender systems, recommending a list of items has also been addressed in cascading bandits [KSWA15]. However, while our objective is to identify the items that each input likes, their focus is on learning the top $K$ items that are liked by most inputs. Another feature of our learning problem is that we can only learn from positive examples. PAC learning for binary classification from positive examples has been studied in the literature [Den98, DDGL99, LDG00, BD20].
Multi-label learning [McC99, SS00] has been an area of study in machine learning, with various, primarily experimental approaches (see, e.g., [EW01, PC11, KVJ12] and [ZZ14, BTDK22] for surveys). In multi-label learning, the training set consists of examples, each associated with multiple labels rather than just one. The goal is to train a model that can learn the relationships between the features of each example and all its labels. At test time, the learner predicts a list of labels for new examples, aiming to capture all the relevant labels, rather than just a single one.
[…]